OpenStudy (anonymous):

If the amount of air in a spherical balloon is 288 cu. in., what is the radius of the balloon?

OpenStudy (wmj259):

Do you know the equation for the Volume of a sphere?

OpenStudy (wmj259):

\[Volume (sphere)= \left( 4\div3 \right) \times \pi \times r ^{3}\]